# Quantifying the dynamic structural resilience of international staple food trade networks: An entropy-based approach

### v2

@@ 6.2 Findings and managerial implications
The results reveal several interesting findings and inspired by these, we summarize several managerial implications.
@@
First, efficiency remains the primary driver of resilience, while current international trade networks tend to exhibit greater redundancy within the $\alpha$ trade-off framework. Furthermore, we find that clearer community partitions can enhance resilience by promoting denser intra-community trade flows and improving overall trade efficiency, as observed in networks for maize, rice, and wheat. At present, therefore, strengthening intra-community trade linkages through regional alliances (e.g., NAFTA, MERCOSUR, G20, ASEAN, and AfCFTA) would greatly contribute to maintaining resilience of trade in these foods (de Raymond et al., 2021), particularly during major disruptions such as the COVID-19 pandemic (Wei et al., 2025). In contrast, the soybean trade network appears more vulnerable due to the highly monopolistic positions of major economies such as the United States, Brazil, and China, resulting in lower modularity and efficiency. Hence, in terms of trade security and economic benefit, economies should strengthen regional production cooperation and technical exchanges to improve soybean production capacity and trade efficiency (Ma et al., 2024). This would not only improve the diversification of trade linkages and risk resistance but also jointly ensure domestic soybean supply and promote industrial development.
@@
Second, rice exhibits distinct trends and characteristics compared to other staple foods, necessitating more targeted attention. We find that several topological properties—such as network size and clustering coefficient—are positively correlated with resilience for maize, soybean, and wheat but not for rice. The resilience of the rice trade network tends to decline with increasing globalization, which may be attributed to its inherent edibility. However, more critical attention should be directed toward the Asia-centered monocentric structure in the rice trade and the frequent export restrictions imposed by major exporters, both of which fundamentally undermine resilience. Hence, expanding production and liberalizing rice exports are essential for improving rice trade resilience. On one hand, expanding rice production outside of Asia is important for enhancing diversification (Puma et al., 2015). This would not only mitigate the risks of supply disruptions caused by events—such as climate extremes, exports bans, and competitive underpricing—in Asia, but also lower rice access costs in other regions and generate income for many developing countries. On the other hand, liberalizing rice exports of major producers, while ensuring sufficient domestic supply, would not only improve overall trade resilience, but also benefit domestic farmers and economy (Valera et al., 2024).
Third, the underlying discourse on food trade has shifted from flow monopolies to strategic interactions within relational networks, placing greater emphasis on the importance of connectivity (Sartori and Schiavo, 2015). Using index decomposition analysis, we find that the concentration effects of trade flows continue to be dominant drivers of changes in resilience with the growing influence of interrelationships among economies in recent years. These effects are further substantiated by the identification of critical contributors to resilience. Economies with moderate trade volumes or low network centrality—such as Paraguay and Kazakhstan—can significantly contribute to system resilience due to their indispensable roles in specific products or strategic geographies. Therefore, for many developing economies, there is an indication that strategic trade positioning may surpass the indiscriminate pursuit of trade volume in enhancing contributions to the international food trade resilience. Moreover, developing and maintaining robust and diversified international food transport systems is essential for improving access to food across various terrains and urban landscapes. Similarly, strengthening food trade infrastructures—such as ports and railways—to ensure international trade connectivity is equally necessary (FAO, 2021).
@@
Fourth, core food trade networks may undermine overall system efficiency, whereas peripheral networks play a critical role in sustaining resilience by enabling risk diversification through alternative trade pathways. Within the $\alpha$ trade-off framework, we find that one of the main factors limiting efficiency in current trade systems is the overconcentration of flows along core trade links, particularly for soybeans. Therefore, guaranteeing food security requires not only strong trade linkages among developed economies but also greater efforts to establish freer trade bridges between developed and undeveloped economies. Such efforts would not only contribute to reducing hunger in developing economies but also help diversify trade risks in developed economies and enhance systemic resilience. As noted by Mujahid and Kalkuhl (2016), all economies are expected to benefit from a free and equitable food trade system that operates without discrimination.
@@ 6.3 Limitations and future research directions
Despite the contributions of our study, several limitations warrant further investigation. First, we do not account for food supply and demand dynamics when adopting the leave-one-out approach. In reality, when an economy or trade relationship is disrupted, affected economies typically adjust their trade strategies to rebalance supply and demand, triggering cascading effects that reshape flow distributions across the entire network. It is therefore important to clarify that our approach is designed solely to assess the contributions of economies or trade links to network resilience. Recent studies have employed agent-based models and stress-testing approaches to capture the dynamics of risk propagation in food systems under shock scenarios (Kuhla et al., 2025; Diem et al., 2025), and have further extended the analytical scope to multilayer cascading risks (Naqvi et al., 2020; Laber et al., 2023). Future research could build on these approaches by simulating trade flow redistributions under various shocks, which would provide deeper insights into the resilience of the international food trade system. Second, we do not apply causal inferences in this paper when investigating the interrelations between resilience and other properties. To address this limitation, future studies could apply more advanced methods—such as machine learning—to better identify and quantify the drivers of resilience within the international food trade network. Third, although the entropy-based framework provides a theoretical metric for resilience, its application should be further validated and adapted in socio-economic networks that differ from ecosystems shaped through natural selection (Kharrazi et al., 2013). Future research could investigate whether further autocatalytic effects can emerge within the socio-economic network structure, enabling the replacement of redundant pathways with more efficient alternatives.
@@ Acknowledgment
This work was partly supported by the National Natural Science Foundation of China (72171083) and the Central Universities’ Program for Building World-Class Universities (Disciplines) and Special Development Guidance—Cultural Heritage and Innovation

### v3

@@ 6.2 Findings and managerial implications
The results reveal several interesting findings and inspired by these, we summarize managerial implications. First, efficiency remains the primary driver of resilience, while current international trade networks tend to exhibit greater redundancy within the $\alpha$ trade-off framework. Our findings align with views of Kharrazi et al. (2013), which suggests that natural evolutionary processes are generally more effective at eliminating excess redundancy than human-designed systems. Furthermore, we find that clearer community partitions can enhance resilience by promoting denser intra-community trade flows and improving overall trade efficiency (see the analysis in Appendix B), as observed in networks for maize, rice, and wheat. Therefore, as de Raymond et al. (2021) note, governments and regional alliances (e.g., NAFTA, MERCOSUR, G20, ASEAN, and AfCFTA) could help to maintain resilience of trade in these foods by fostering denser intra-community trade flows. For example, developing common storage facilities, harmonizing trade standards, and reducing non-tariff barriers within regional blocs would strengthen intra-community linkages, particularly during major disruptions such as the COVID-19 pandemic (Wei et al., 2025). In contrast, the soybean trade network appears more vulnerable due to the imbalance between efficiency and redundancy, largely attributable to the highly monopolistic positions of major economies such as the United States, Brazil, and China, resulting in lower modularity and efficiency (Li et al., 2024). Hence, in terms of trade security and economic benefit, governments of soybean-importing economies could invest in joint regional production programs, support farmer cooperatives, and strengthen technology transfer agreements that enhance local production capacity and efficiency (Ma et al., 2024). This would not only diversify trade linkages, enhance risk resistance, and jointly ensure domestic soybean supply while promoting industrial development.
@@
Second, rice exhibits distinct trends and characteristics compared to other staple foods, necessitating more targeted attention (Burkholz and Schweitzer, 2019). We find that several topological indicators—such as network size and clustering coefficient—are positively correlated with resilience for others but not for rice. The resilience of the rice trade network tends to decline with increasing globalization, which may be attributed to its inherent edibility. However, more critical attention should be directed toward the Asia-centered monocentric structure in the rice trade and the frequent export restrictions imposed by major exporters (Davis et al., 2021), both of which fundamentally undermine resilience. For example, Baum et al. (2025) have confirmed that Indian rice export ban substantially alter global food availability. Hence, expanding production and liberalizing rice exports are essential for improving rice trade resilience. On one hand, international organizations (e.g., FAO, WTO) should encourage diversification of rice production beyond Asia through targeted investment projects and development aid (Puma et al., 2015). This would not only mitigate the risks of supply disruptions caused by events—such as climate extremes, exports bans, and competitive underpricing—in Asia, but also lower rice access costs in other regions and generate income for many developing countries. On the other hand, governments of major exporting countries could design export policies that allow for controlled liberalization during normal times while ensuring sufficient domestic supply. As Valera et al. (2024) note, liberalizing rice exports of major producers, while ensuring sufficient domestic supply, would benefit domestic farmers and economy.
Third, the underlying discourse on food trade has shifted from flow monopolies to strategic interactions within relational networks, placing greater emphasis on the importance of connectivity (Sartori and Schiavo, 2015). Using index decomposition analysis, we find that the concentration effects of trade flows continue to be dominant drivers of changes in resilience, while the influence of interrelationships among economies is growing in recent years. These effects are further substantiated by the identification of critical contributors to resilience. Although our findings are broadly consistent with several studies in confirming the systemic importance of large exporters including the United States, Brazil, India, and Russia (Kuhla et al., 2025; Laber et al., 2023; Baum et al., 2025), the leave-one-out analysis reveals that some peripheral economies can exert disproportionate influence on the overall resilience by sustaining irreplaceable trade pathways, which complements the above findings. For example, the existence of the link between Paraguay and Brazil improves Paraguay’s indispensable role in networks for maize and rice, while Kazakhstan secures a critical position as a crisis replacement corridor. This implies that developing economies may enhance their contributions to resilience not only by increasing volumes but by adopting strategic trade positioning, such as specializing in niche commodities or maintaining bilateral agreements with key partners. For policymakers in developing countries, strengthening transport and storage infrastructure (e.g., ports, railways, grain silos) is crucial for realizing such strategies (FAO, 2021). For donor agencies and international lenders, providing financial support for infrastructure projects directly improves resilience by enabling diversified trade routes and better market access.
@@
Fourth, core food trade networks may undermine overall system efficiency due to the overconcentration of flows, whereas peripheral networks play a critical role in sustaining resilience by enabling risk diversification through alternative trade pathways. Within the $\alpha$ trade-off framework, we find that one of the main factors limiting efficiency in current trade systems is the overconcentration of flows along core trade links, which aligns with previous findings (Karakoc and Konar, 2021), particularly for soybean. Developed economies could therefore complement their strong intra-core connections with policies that actively expand trade with developing economies, such as reducing tariffs on staple imports or supporting trade facilitation programs. This would simultaneously reduce hunger risks in vulnerable economies and diversify risks for developed ones. International organizations should play a coordinating role by promoting transparent and equitable trade rules, ensuring that peripheral actors are integrated into the global food trade system. As noted by Mujahid and Kalkuhl (2016), all economies are expected to benefit from a free and equitable food trade system that operates without discrimination.
@@ 6.3 Limitations and future research directions
Despite the contributions of our study, several limitations warrant further investigation. First, although the entropy-based framework provides a theoretical metric for resilience, its application should be further validated and adapted in socio-economic networks that differ from ecosystems shaped through natural selection (Kharrazi et al., 2013). Also, the trade-off is relative in this framework, so optimal resilience may occur even when both efficiency and redundancy are low. Therefore, the trends in efficiency and redundancy need to be jointly considered when assessing system resilience. Future research could investigate whether further autocatalytic effects emerge within the socio-economic network structure, thereby enabling redundant pathways to be replaced by more efficient alternatives. Moreover, building on different concepts and frameworks, our inferences about resilience differ from those in some studies. For example, we argue that the directed ring network is overly efficient and non-resilient (see Appendix A4), whereas Karakoc and Konar (2021) regard it as resilient. Hence, discussions toward a unified concept and framework of resilience warrant further development.
Second, we do not consider additional external drivers of resilience as other studies, such as geopolitical conflicts (Laber et al., 2023), extreme climate events (Davis et al., 2021), and trade policy uncertainty (Li et al., 2024). Moreover, we do not apply causal inferences in this paper when investigating the interrelations between resilience and other drivers. To address these limitations, future studies could apply more advanced methods—such as machine learning—to better identify and quantify the drivers of resilience within the international food trade network, providing a more comprehensive understanding of resilience dynamics. Also, addressing these limitations could provide more robust tools for policymakers and practitioners in managing trade systems and making them more resilient.
@@
Third, we do not account for food supply and demand dynamics when adopting the leave-one-out approach. In reality, when an economy or trade relationship is disrupted, affected economies typically adjust their trade strategies to rebalance supply and demand, triggering cascading effects that reshape flow distributions across the entire network. It is therefore important to clarify that our approach is designed solely to assess the contributions of economies or trade links to network resilience. Recent studies have employed agent-based models and stress-testing approaches to capture the dynamics of risk propagation in food systems under shock scenarios (Kuhla et al., 2025; Diem et al., 2025), and have further extended the analytical scope to multilayer cascading risks (Naqvi et al., 2020; Laber et al., 2023). Future research could build on these approaches by simulating trade flow redistributions under various shocks, which would provide deeper insights into the resilience of the international food trade system.
7 Conclusion
Based on the FAO dataset from 1986 to 2022, we construct international trade networks for maize, rice, soybean, and wheat, where economies are represented as nodes and trade relationships among them as directed and weighted links. To address the need of resilience measurement in the networked system, we employ an entropy-based approach grounded in information theory. Subsequently, we explore the correlations between resilience and topological indicators. We also employ the index decomposition analysis to quantify the relative contributions of several properties to resilience, including efficiency, redundancy, trade flows concentration, and interrelationship among economies. Furthermore, we apply the leave-one-out approach to identify critical economies and trade relationships from a resilience perspective. Given the pronounced core-periphery structure of networks, we also discuss the roles of core and peripheral components in sustaining network resilience.
Our findings provide an empirical evidence for an emerging consensus on balancing multiple principles of resilience. Given that current international trade networks are more redundant within the $\alpha$ trade-off framework, enhancing modularity and efficiency to improve resilience represents the next step in optimizing trade networks. Moreover, our findings not only suggest the dominant role of trade flow concentration, but also highlight the growing influence of interrelationships among economies, which further emphasizes the indispensable roles of some peripheral economies and trade links in sustaining resilience. From this perspective, our research highlights the attention to peripheral economies and trade relationships, while advocating for strengthened intra-regional food trade and more liberalized cross-regional food trade, thereby contributing to global food security.
@@ Acknowledgment
This work was partly supported by the National Natural Science Foundation of China under Grant 72171083 and the Central Universities’ Program for Building World-Class Universities (Disciplines) and Special Development Guidance—Cultural Heritage and Innovation.
@@
Appendix
Appendix A Special cases among different networks
A1 All volume on one link
Suppose that $f_{kl}=s$ for link $\varepsilon_{kl}$ from node $k$ to node $l$ and $f_{ij}=0$ for all other links. In this case, we have $p_{kl}=1$, $p_{k}^{\mathrm{out}}=1$, $p_{l}^{\mathrm{in}}=1$, and $p_{ij}=p_{i}^{\mathrm{out}}=p_{j}^{\mathrm{in}}=0$ for $\varepsilon_{ij}\neq\varepsilon_{kl}$. The joint entropy $H$ can be calculated as
$H=-p_{kl}\ln p_{kl}-\sum_{\begin{subarray}{c}(i,j)\neq(k,l)\end{subarray}}p_{ij}\ln p_{ij}=-\underbrace{1\cdot\ln(1)}_{\text{term }(k,l)}-\sum_{\begin{subarray}{c}(i,j)\neq(k,l)\end{subarray}}\underbrace{0\cdot\ln(\cdot)}_{=0}=0,$
while $e$ and $r$ can be calculated as
$e=\underbrace{-1\cdot\ln\frac{1}{1\cdot 1}}_{\text{term }(k,l)}-\sum_{\begin{subarray}{c}(i,j)\neq(k,l)\end{subarray}}\underbrace{0\cdot\ln(\cdot)}_{=0}=0$
and
$r=\underbrace{1\cdot\ln\frac{1}{1}+1\cdot\ln\frac{1}{1}}_{\text{term}(k,l)}-\sum_{\begin{subarray}{c}(i,j)\neq(k,l)\end{subarray}}\underbrace{0\cdot\ln(\cdot)}_{=0}=0,$
which indicates that the system is deterministic (i.e., has no uncertainty).
A2 No volume on some link
Suppose that link $\varepsilon_{ij}$ is structurally present, but $f_{ij}=p_{ij}\to 0$. Four situations may arise:
  1. 1)
    Node $i$ has no other outgoing links and node $j$ has no other incoming links such that $p_{i}^{\mathrm{out}}=p_{j}^{\mathrm{in}}=p_{ij}\to 0$;
  2. 2)
    Node $i$ has no other outgoing links but node $j$ has other incoming links such that $p_{i}^{\mathrm{out}}=p_{ij}\to 0$ but $p_{j}^{\mathrm{in}}\neq 0$;
  3. 3)
    Node $i$ has other outgoing links but node $j$ has no other incoming links such that $p_{j}^{\mathrm{in}}=p_{ij}\to 0$ but $p_{i}^{\mathrm{out}}\neq 0$;
  4. 4)
    Nodes $i$ has other outgoing links and node $j$ has other incoming links such that $p_{i}^{\mathrm{out}}\neq 0$ and $p_{j}^{\mathrm{in}}\neq 0$.
In all of these cases, the mutual information between $i$ and $j$ can still be calculated as
$\lim_{p_{ij}\to 0^{+}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=0\cdot\ln(\cdot)=0,$
and the conditional entropy between $i$ and $j$ can still be calculated as
$\lim_{p_{ij}\to 0^{+}}\left(p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}}+p_{ij}\ln\frac{p_{ij}}{p_{j}^{\mathrm{in}}}\right)=0\cdot\ln(\cdot)=0,$
which indicates that zero-flow pairs do not contribute to the total efficiency and redundancy.
A3 Directed fully connected trade network
Consider a directed fully connected trade network, i.e., a pair of reciprocal links exists between any two nodes. For simplicity, we suppose that $f_{ij}=p_{ij}=1/N_{\mathcal{E}}$, where $N_{\mathcal{E}}=N_{\mathcal{V}}(N_{\mathcal{V}}-1)$. Hence, the efficiency and redundancy can be calculated as
$e=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{1}{N_{\mathcal{V}}(N_{\mathcal{V}}-1)}\ln\frac{N_{\mathcal{V}}(N_{\mathcal{V}}-1)}{(N_{\mathcal{V}}-1)^{2}}=\ln\frac{N_{\mathcal{V}}}{N_{\mathcal{V}}-1},$
and
$r=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}^{2}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{1}{N_{\mathcal{V}}(N_{\mathcal{V}}-1)}\ln(N_{\mathcal{V}}-1)^{2}=2\ln(N_{\mathcal{V}}-1),$
where $r$ exhibits a logarithmic increase but $e$ converges to 0 if $N_{\mathcal{V}}$ is large.
A4 Directed ring trade network
Consider a directed ring trade network, i.e., each node has exactly one in-flow pathway and one out-flow pathway. For simplicity, we suppose that $f_{ij}=p_{ij}=1/N_{\mathcal{E}}$, where $N_{\mathcal{E}}=N_{\mathcal{V}}$. Hence, the efficiency and redundancy can be calculated as
$e=\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{1}{N_{\mathcal{V}}}\ln\frac{N_{\mathcal{V}}^{2}}{N_{\mathcal{V}}}=\ln N_{\mathcal{V}},$
and
$r=-\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}p_{ij}\ln\frac{p_{ij}^{2}}{p_{i}^{\mathrm{out}}p_{j}^{\mathrm{in}}}=\sum_{i=1}^{N_{\mathcal{V}}}\sum_{j=1}^{N_{\mathcal{V}}}\frac{1}{N_{\mathcal{V}}}\ln\frac{N_{\mathcal{V}}^{2}}{N_{\mathcal{V}}^{2}}=0,$
where $e$ exhibits a logarithmic increase if $N_{\mathcal{V}}$ is large. It also indicates that the system is deterministic (i.e., has no uncertainty).
These cases confirm that the $\alpha$ resilience framework is well-defined for non-negative flows ($f_{ij}\geq 0$ and $p_{ij}\geq 0$), and behaves consistently in scenarios such as deterministic cases, zero-flow systems, fully connected networks, and ring networks.
Appendix B Modularity and resilience
Modularity can be expressed as:
$Q=\frac{1}{N_{\mathcal{V}}\langle s\rangle}\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\left(f_{ij}-\frac{s_{i}s_{j}}{N_{\mathcal{V}}\langle s\rangle}\right)\delta(c_{i},c_{j})=\frac{1}{s}\sum\limits_{i=1}^{N_{\mathcal{V}}}\sum\limits_{j=1}^{N_{\mathcal{V}}}\left(f_{ij}-\frac{s_{i}s_{j}}{s}\right)\delta(c_{i},c_{j}),$
where $s$ denotes the the total system flow. Relatively large modularity values between 0.3 and 0.7 are considered an indicator of a well-structured network partition (Newman, 2004), which represents that the actual flows are higher than the expected flows within communities, expressed as:
$f_{i,j}\gg\frac{s_{i}s_{j}}{s},$
which denotes that flows within communities are inter-dependent, corresponding to the higher efficiency in Eq. (1) and pmi in Eq. (14). Since the current networks are more redundant within the $\alpha$ framework, the higher efficiency is, the higher resilience is. Hence, for the current international food trade networks with stronger community structures, $Q\propto e\propto R$.
Appendix C Community partition for food trade networks
In this paper, we use the Louvain algorithm for community detection and examine the evolution of community structures across three periods (Fig. C11). Economies within the same community are typically clustered in contiguous geographic regions, a pattern that can be largely explained by transport costs, regional trade policies, and geopolitical ties. For example, European countries consistently form a single community; North Africa and the Middle East often constitute a distinct bloc or alternatively maintain stable linkages with Europe; and South American countries largely cluster together. Geographic proximity fosters closer trade linkages among neighboring economies, thereby facilitating risk distribution at lower transportation costs in the event of localized shocks and ultimately enhancing overall resilience.
Beyond geography, food-specific attributes also shape community evolution. For soybean, the persistent dominance of the United States, Brazil, and Argentina as major producers and exporters has fostered inter-community connectivity but simultaneously constrained the emergence of diversified community structures. This concentration weakens modular boundaries and reduces the contribution of community diversity to resilience. By contrast, rice and wheat production and exports are more geographically dispersed, resulting in more distinct and diverse community structures, thereby sustaining resilience by strengthening community-based risk buffering.
Figure C11: Community structures in the international trade networks for maize, rice, soybean, and wheat. In this figure, a blank indicates that: 1) no economy here, 2) the economy did not participate in trade, or 3) data are missing. a-d, In 1998. c-h, In 2010. i-l, In 2022.